OpenStudy (julierosehead):

Input (x): 0, 1, 2, 4 Output (y): 4, 3, 2, 0 Write a rule for this function.. Thank you: Note: i have a feeling that the answer is "there is no rule because it is not a function" but just making sure. thanks

OpenStudy (karim728):

what makes u think its not a function?

563blackghost (563blackghost):

`A function correlates with each element of a set with only one element of another set.` Since there is one y-value for every of x this is a function.

OpenStudy (julierosehead):

I don't know, I didn't quite understand this part of algebra, but i really would like to know the rule for this function, i usually do very well in math, this is the last question i have in math

563blackghost (563blackghost):

To find the function rule we can determine the slope of two points. \(\huge\bf{\frac{rise}{run} \rightarrow \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{3-4}{1-0}=-1}\) We are also given at when the line passes through the y-axis which is `4` so the y-intercept is `4`. So the function rule is. \(\huge\bf{y=-x+4}\)

563blackghost (563blackghost):

@julierosehead

OpenStudy (julierosehead):

thank you so much!!

563blackghost (563blackghost):

No problem :)